Roger has $3.15 in quarters dimes and nickels he has twice as many quarters as dimes and two more dimes than nickels. how many of each kind are there?
Answer by FrankM(1040)

5N+10D+25Q=315 << We mutiply the number of Nickels (N) by its value 5, to get the value of all nickels, 10 for dimes, 25 for quarters.
Q=2D
D=N+2
Now, 3 equations, in 3 unknowns.
Since Q=2D, Q=2(N+2)
5N+10(N+2)+25(2(N+2))=315
5N+10N+20+50N+100=315
65N+120=315
65N=195
N=3, D=5, Q=10
Check -
5*3+10*5+25*10=315
15+50+250 = 315